OpenStudy (anonymous):

a box is to be made by cutting out the corners of a square piece of cardboard and folding the edges up. If 3 inch squares are to be cut out of the corners and the box contains 243 cubic inches, what is the length of a side of the original cardboard square?

OpenStudy (anonymous):

Let us assume side of initial square be x. Now, after cutting squares of 3in from the corners and making a box by folding the sides, will form a box of volume = (x-6)(x-6)(3). So, (x-6)(x-6)(3)=243 (x-6)(x-6)=81 x-6=9 x=15in.